OpenStudy (anonymous):

Hello! what`s the integral of ((2x+1)/x(x-2))dx?

OpenStudy (anonymous):

Ok so i went down the wrong path first time but i have it now, have you thought about splitting this integral into partial fractions?

OpenStudy (anonymous):

ok i`m really suck at that and what stoping me now is that in my examples we made it to look like this : \[\int\limits \frac{ 2 }{ x^2-1 } = \frac{ 2 }{ (x+1) (x-1) }\] from that i can do it myself but i`m really slow now :D and it`s been a wile since i did math last time, so i`m stuck on making bottom part of a fraction to have 2 ( ) going to google meanwile partial fractions

OpenStudy (anonymous):

unless... because top part is higher then bottom and my formula is not applicable here...

OpenStudy (anonymous):

ok so in the example you have given me above to split this into partial fractions this is the first step:\[\int\limits_{}^{}\frac{ 2 }{ x^{2}-1 } = \int\limits_{}^{}\frac{ 2 }{ (x+1)(x-1) } = \int\limits_{}^{}\frac{ A}{ (x+1) }\times \frac{ B }{ (x-1) }\] So now you need to calculate A and B Do you know/remeber how do do that? If now just say an I can explain

hartnn (hartnn):

\(\huge \dfrac {2x+1}{x(x-2)}=\dfrac{A}{x}+\dfrac{B}{x-2}\) ^partial fractions

hartnn (hartnn):

that would give you \(\Large 2x+1 = A(x-2)+Bx \) to find A, plug in x =0 to find B, plug in x=2 what do you get ?

OpenStudy (anonymous):

yes from that i can work with.. will it be like that : \[2=A*(x-1)+B*(x+1)\] x=-1 2=A*(-2)+0 A=2/-2= -1 x=1 2=0+B*2 B=1 \[\int\limits (\frac{ -1 }{ x+1 }) + (\frac{ 1 }{ x-1 })dx=-\int\limits \frac{ 1 }{x+1 }+\int\limits \frac{ 1 }{ x-1 }dx=-\ln \left| x+1 \right| +\ln \left| x-1 \right| + R\] what confusing me there is that how do i get can i make my original question to look like A/x + B/x-2 and go from there.. or do i have to have ()() two of thouse at the bottom?

OpenStudy (anonymous):

aha.. so i do can go that way.. give me a sec

hartnn (hartnn):

sure, take your time :) when you get A,B, tell me, so that i can verify :)

OpenStudy (anonymous):

x=0 0+1=A*(-2) A=-(1/2) x=2 4+1=0 + B*2 B=5/2 so far..

hartnn (hartnn):

correct! :)

hartnn (hartnn):

\(\Huge \int \dfrac{-1}{2x}dx+\int \dfrac{5/2}{x-2}dx\) can you integrate/proceed ? ask if u get stuck...

OpenStudy (anonymous):

then next it will look like this? : \[\int\limits \frac{ 2x+1 }{ x(x-2) }=\int\limits (\frac{ -\frac{ 1 }{ 2 } }{ x }+\frac{ \frac{ 5 }{ 2 } }{ x-1 })dx=-\int\limits \frac{ \frac{ 1 }{ 2 } }{ x }dx+\int\limits \frac{ \frac{ 5 }{ 2 } }{ x-1 }dx\]

hartnn (hartnn):

yes, good!

OpenStudy (anonymous):

ow yeah.. x-2.. misread.. so far :D but this part gonna be tricky i think.. give me a moment

OpenStudy (anonymous):

-ln|2x| + 5/2*ln|x-1| <--- is this final answer?

hartnn (hartnn):

absolutely correct! :) just add +c in the end :)

OpenStudy (anonymous):

ow yeah.. damn thank you veeery much.. now i just need to do more exercises to fix it in my brain... thanks a lot.. do i give you some likes here or feedbacks to increase your rating or "thank you" is enough here?

hartnn (hartnn):

i am just happy that you understood, and hope that you could solve similar problem on your own. That would suffice or me :) and since you're new here, let me take this opportunity to welcome you, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

thank you once more :D gonna try to do another one meanwhile, about surface area and volume... but i`ll probably gonna need some help on that too.. is it ok to continue this topic or better to create a new one?

hartnn (hartnn):

it'd better i you 'Close' this post and ask new question in new post. because i may not be online to help you, but there are many others here willing to help :)